Write your explanations to the right of each step as you would explain the following
algebraic problem to a learner:
2.3x=162
3x= 81
3x=34
x = 4
1
Expert's answer
2020-09-17T16:57:39-0400

We are given "2.3^x = 162"

The equation can be written clearly as "2(3^x) = 162"


Step 1:

Divide both sides of the equation by 2 to reduce it.

"=> \dfrac {2(3^x)}{2} = \dfrac {162}{2}"


"=> 3^x = 81"


Step 2:

Write 81 to base 3 such that the expressions on both the L.H.S and R.H.S are raised to the same base, 3.


"=> 3^x = 3^4"


Step 3:

Equate powers since both the L.H.S and the L.H.S of the equation are raised to the same base.

"Note: \space a^m = a^n => m = n"

"\therefore \space 3^x = 3^4 => x = 4"


Step 4:

Conclude,

"If \space 2.3^x = 162 , \space \space Then \space x= 4"